If a lender makes a simple loan of $800 for 2 years and charges 3%, then the amount that the lender receive at maturity is $?. Round your response to the nearest two decimal place

If a lender makes a simple loan of $2500 for one year and charges $30 interest, then the simple interest rate on that loan is ?% Round your response to the nearest whole number.

If a borrower must repay $106.50 one year from today in order to receive a simple loan of $100 today, the simple interest on this loan is?

A.6.5%

B.5.0%

C.65%

D.6.0%
Answer #1

1. Ans: $848.72

Explanation:

Maturity value or Future value ( F ) = P * ( 1 + i )n

= $800 ( 1 + 0.03 )2

= $800 ( 1.03 )2

= $800 * 1.0609

= $848.72

2. Ans: 1.2 %

Explanation:

Total interest earned in 1 year = $30

Total loan amount = $2500

So the simple interest on that loan = (30 / 2500 ) *100

= 0.012 * 100

= 1.2

3. Ans: A ) 6.5%

Explanation:

Total interest earned in 1 year = $106.50 - $100 = $6.50

So the simple interest on this loan = (6.50 / 100 )* 100

= 0.065 * 100

= 6.5